# On Purely Private Covariance Estimation

### v1

@@ Abstract
We present a simple perturbation mechanism for the release of $d$-dimensional covariance matrices $\Sigma$ under pure differential privacy. For large datasets with at least $n\geqslant d^{2}/\varepsilon$ elements, our mechanism recovers the provably optimal Frobenius norm error guarantees of [Nik23], while simultaneously achieving best known error for all other $p$-Schatten norms, with $p\in[1,\infty]$. Our error is information-theoretically optimal for all $p\geqslant 2$, in particular, our mechanism is the first purely private covariance estimator that achieves optimal error in spectral norm.
For small datasets $n<d^{2}/\varepsilon$, we further show that by projecting the output onto the nuclear norm ball of appropriate radius, our algorithm achieves the optimal Frobenius norm error $O(\sqrt{d\;\text{Tr}(\Sigma)/n})$, improving over the known bounds of $O(\sqrt{d/n})$ of [Nik23] and ${O}\big(d^{3/4}\sqrt{\text{Tr}(\Sigma)/n}\big)$ of [DLY22].
@@ 1 Introduction
Because of its widespread use in training prediction models, the task of releasing a private version of $\Sigma$ has received significant attention (see [ADK+19, DLY22, Nik23, CAdE+24] and references therein). In contrast to ad-hoc private training methods, the advantage of working with a synthetic version $\hat{\Sigma}$ of the covariance is that one can replace $\Sigma$ with $\hat{\Sigma}$ and directly run standard training algorithms. This approach not only fits seamlessly in existing training pipelines, but typically also leads to lower privacy costs, as these are paid only once, upon computing the private estimate.
The de facto standard for privacy is differential privacy ([DMNS06]). A randomized algorithm $\mathcal{A}$ is said to be $(\varepsilon,\delta)$-differentially private if, for any pair of datasets $X,X^{\prime}\subseteq\mathbb{R}^{d}$ differing in at most $1$ element, and for any measurable event $S$ in the range of $\mathcal{A},$ it holds
@@
The setting $\delta=0$ is usually refer to as pure privacy, while the setting $\delta>0$ is called approximate differential privacy. Pure privacy is the strictest notion of privacy, it disallows any complete privacy breach and naturally offers privacy protection to groups of elements in the datasets, as opposed to just individuals (see [Nik23] for a more comprehensive discussion).
In the context of private covariance estimation, the difference between pure and approximate differential privacy is surprisingly stark. For $\delta>0,$ efficient algorithms can achieve accuracy $\lVert\hat{\Sigma}-\Sigma\rVert_{\mathrm{F}}^{2}\leqslant O\left(\sqrt{\log(1/\delta)}/\varepsilon\right)\cdot\min\left\{d/n,\,d^{1/4}\sqrt{\varepsilon/n}\right\}$ and this bound is known to be tight ([NTZ13, DNT15, KRSU10]). In contrast, pure privacy lower bounds for the special case of $2$-way marginals already rule out a Frobenius-norm error smaller than $\min\left\{d^{3/2}/\varepsilon n,\sqrt{d/n\varepsilon}\right\}$ ([HT10, Nik23]). Here the second term is smaller for small sample sets $n<d^{2}/\varepsilon$.
[Nik23] introduced a variant of the projection mechanism, based on the Johnson-Lindestrauss transform ([Joh84]) matching these bounds up to constant factors. For datasets with points of small average Euclidean norm, [DLY22] designed a pure differentially private algorithm achieving a Frobenius norm error of $\tilde{O}(d^{3/4}\sqrt{\operatorname{Tr}\Sigma/n}+\sqrt{d}/n),$ thus going beyond [Nik23] for the special case of matrices satisfying $\operatorname{Tr}\Sigma\ll O(1/\varepsilon)\cdot\min\left\{1/n,\,d^{-1/4}\right\}.$
@@ 1.1 Results
Efficient and streamlined samplers for the nuclear-Laplace law are easy to construct and can be seen as a specific instantiation of the $K$-norm mechanism of [HT10]. Therefore the algorithm runs in polynomial time.
@@
Theorem 1.1 guarantees bounded error simultaneously for all Schatten norms, including the nuclear norm ($p=1$) and the spectral norm ($p=\infty$), implying spectral discrepancies between the true covariance and the pure differentially private estimate are small under any natural reweighing. For large datasets $n\geqslant d^{2}/\varepsilon$ –typical of real-world applications– the Theorem recovers the guarantees of [Nik23] for the Frobenius norm error. Furthermore, as for $p\geqslant 2$
@@
the error of Theorem 1.1 is optimal for $p\geqslant 2$ in the large sample regime $n\geqslant d^{2}/\varepsilon$. Indeed, the information-theoretic lower bound ([Nik23]) shows that if $n\geqslant d^{2}/\varepsilon$, no pure DP estimator can achieve error $o(d^{3/2}/\varepsilon n)$ in Frobenius norm, and hence no estimator can achieve error $o(d^{1+1/p}/\varepsilon n)$ in Schatten $p$-norm for $p\geqslant 2$. In particular, our spectral norm bound $O(d/\varepsilon n)$ is information-theoretically optimal in the large sample regime $n\geqslant d^{2}/\varepsilon$.
@@
Theorem 1.2 always recovers the result of [Nik23], but improves over it whenever $\operatorname{Tr}\Sigma<d^{2}/n.$ As values of the trace of the covariance matrix can be significantly smaller than $1$, Theorem 1.2 can lead to important accuracy improvements both in the small sample regime $n\leqslant d^{2}/\varepsilon$ where the improvement is of the order $O(1/\operatorname{Tr}\Sigma)$, and in the large sample regime. Theorem 1.2 also strictly improves over the result of [DLY22] (that also captured the dependence on $\operatorname{Tr}(\Sigma)$) in all regimes by a $d^{1/4}$ factor. Finally, we remark that the term $\frac{\sqrt{d\log d}}{\varepsilon n}$ is unavoidable ([KLSU19]).
@@ 2 Techniques
To illustrate our ideas we start by reviewing the algorithm of [Nik23]. On a high level, this consists of three steps: (1) project the data onto a low dimensional space, (2) add noise via a $K$-norm mechanism, and (3) project the result back onto the space of covariance matrices (the intersection of the nuclear unit ball and the positive semidefinite cone). As the cost of traditional perturbation mechanisms tends to be proportional to the ambient dimension, the goal of the first step is to reduce this privacy cost by embedding the dataset into a small subspace, thus paying a cost proportional only to the effective dimension of the data. The result $\tilde{\Sigma}$ of steps (1) and (2) can then be shown to be private and close to (the projection of) $\Sigma$ up to a Frobenius distance of order $O(d^{3/2}/n\varepsilon)$. The third step is an orthogonal projection onto a convex set containing the ground truth, thus it cannot increase the $\ell_{2}$-distance to $\Sigma$, but, remarkably, for small datasets can provably shrink it to $O(\sqrt{d/n\varepsilon}).$
Despite its utility in terms of the Frobenius norm error, this algorithm cannot be expected to be accurate in other Schatten norms, in a strong sense. The dimensionality reduction step rely on a Johnson-Lindestrauss transform, but such mappings are known not to exist for spaces that are far from being Euclidean and, more specifically, for $p$-norm space with $p\in[1,\infty]\setminus\{2\}$ ([BC05, LMN05, JN10]).
@@
To bypass this inherent limitation, we do not attempt to explicitly embed the dataset in a low-dimensional space and instead carefully craft the additive noise to inject. Since adjacent covariance matrices $\Sigma,\Sigma^{\prime}$ satisfy $\lVert\Sigma-\Sigma^{\prime}\rVert_{*}\leqslant 2/n$, the $K$-norm mechanism ([HT10]) with density $p(Z)$ proportional to $\exp(-\varepsilon n\lVert Z\rVert_{*}/2)$ is $\varepsilon$-DP. Note that it is possible to sample from such distribution efficiently, leveraging any efficient weak separation oracle for the nuclear norm [^3] See Lemma A.2. in [HT10] for more details.. Hence it only remains to bound Schatten norms of $Z$ sampled from this distribution, and in fact, it is enough to bound $\lVert Z\rVert$ by $O(\frac{d}{\varepsilon n})$ since $\lVert Z\rVert_{\textnormal{S}_{p}}\leqslant\lVert Z\rVert\cdot d^{1/p}$.
@@
as [^4] This factorization is standard, and can be found, in particular, in Lemma 1.5.3 from [Chi03].
@@
This implies that to show the desired concentration of the spectral norm, it is enough to study the product of two independent random variables: $R$ and $\max_{i\in[d]}w_{i}$. It is well-known that the distribution of $R$ is [^5] See, for example, Remark 4.2 in [HT10]. $\mathrm{Gamma}(d^{2},\rho)$, and it is well-concentrated around $\rho d^{2}$ (recall that $\rho=2/(\varepsilon n)$). The result would then follow if we could show that with high probability $\max_{i\in[d]}w_{i}\leqslant 1/d+\tilde{O}(1/d^{3/2})$ as by union bound $\left\lVert Z\right\rVert\leqslant O(d^{2}/\varepsilon\cdot n)\cdot O(1/d)\leqslant O(\tfrac{d}{\varepsilon n}).$
@@
Furthermore, a standard strong convexity argument (e.g. the one used in Lemma 5.2 in [CAdE+24]) implies that if $r\geqslant\operatorname{Tr}\Sigma$, then the orthogonal projection $\Pi(\Sigma+Z)$ satisfies
@@
where in the before-last step we used the fact that the spectral norm of $Z$ concentrates around $\Theta(d/n\varepsilon)$ with high probability, and in the last step the radius of the nuclear norm ball. The naive choice $r=1$ would recover the result of [Nik23]. However, this choice can be wasteful. We instead use an $\varepsilon$-DP Laplace estimator of $\operatorname{Tr}\Sigma$. Note that with probability at least $1-\beta$, $r\leqslant O\left(\operatorname{Tr}(\Sigma)+\log(1/\beta)/(n\varepsilon)\right)$, and we get the bound as in Theorem 1.2.
@@ 3 Future work
It is natural to ask whether a generalization of both Theorem 1.1 and Theorem 1.2 is possible. Concretely, it would be interesting to obtain an algorithm such that with high probability
@@
In addition, it would be interesting to see complementary lower bounds. In particular, currently even in the large sample regime $n\geqslant d^{2}/\varepsilon$ it is only known that our result is optimal for $p\geqslant 2$, but it is not known for $p<2$.
@@ 4 Analysis of the additive mechanism
(Remark 4.2 in [HT10]) Let $Z$ have density $p_{\rho}(Z)\propto\exp\left(-\lVert Z\rVert_{*}/\rho\right)$. Then
@@
for a constant $C_{d}>0$, where $d\sigma$ is Lebesgue measure on $\mathbb{R}^{d}$ restricted to $\sigma_{i}\geqslant 0$ and $d\mu_{\mathrm{Haar}}$ denotes Haar probability measure on $\mathrm{O}(d)$ (see Lemma 1.5.3 from [Chi03]).
@@
By , for each LL–Lipschitz $F$ and each $t>0$,

### v2

@@ Abstract
We present a simple perturbation mechanism for the release of $d$-dimensional covariance matrices $\Sigma$ under pure differential privacy. For large datasets with at least $n\geqslant d^{2}/\varepsilon$ elements, our mechanism recovers the provably optimal Frobenius norm error guarantees of [14], while simultaneously achieving best known error for all other $p$-Schatten norms, with $p\in[1,\infty]$. Our error is information-theoretically optimal for all $p\geqslant 2$, in particular, our mechanism is the first purely private covariance estimator that achieves optimal error in spectral norm.
For small datasets $n<d^{2}/\varepsilon$, we further show that by projecting the output onto the nuclear norm ball of appropriate radius, our algorithm achieves the optimal Frobenius norm error $O(\sqrt{d\;\text{Tr}(\Sigma)/n})$, improving over the known bounds of $O(\sqrt{d/n})$ of [14] and ${O}\big(d^{3/4}\sqrt{\text{Tr}(\Sigma)/n}\big)$ of [5].
@@ 1 Introduction
Because of its widespread use in training prediction models, the task of releasing a private version of $\Sigma$ has received significant attention (see [1, 5, 14, 4] and references therein). In contrast to ad-hoc private training methods, the advantage of working with a synthetic version $\hat{\Sigma}$ of the covariance is that one can replace $\Sigma$ with $\hat{\Sigma}$ and directly run standard training algorithms. This approach not only fits seamlessly in existing training pipelines, but typically also leads to lower privacy costs, as these are paid only once, upon computing the private estimate.
The de facto standard for privacy is differential privacy ([6]). A randomized algorithm $\mathcal{A}$ is said to be $(\varepsilon,\delta)$-differentially private if, for any pair of datasets $X,X^{\prime}\subseteq\mathbb{R}^{d}$ differing in at most $1$ element, and for any measurable event $S$ in the range of $\mathcal{A},$ it holds
@@
The setting $\delta=0$ is usually refer to as pure privacy, while the setting $\delta>0$ is called approximate differential privacy. Pure privacy is the strictest notion of privacy, it disallows any complete privacy breach and naturally offers privacy protection to groups of elements in the datasets, as opposed to just individuals (see [14] for a more comprehensive discussion).
In the context of private covariance estimation, the difference between pure and approximate differential privacy is surprisingly stark. For $\delta>0,$ efficient algorithms can achieve accuracy $\lVert\hat{\Sigma}-\Sigma\rVert_{\mathrm{F}}^{2}\leqslant O\left(\sqrt{\log(1/\delta)}/\varepsilon\right)\cdot\min\left\{d/n,\,d^{1/4}\sqrt{\varepsilon/n}\right\}$ and this bound is known to be tight ([13, 7, 11]). In contrast, pure privacy lower bounds for the special case of $2$-way marginals already rule out a Frobenius-norm error smaller than $\min\left\{d^{3/2}/\varepsilon n,\sqrt{d/n\varepsilon}\right\}$ ([8, 14]). Here the second term is smaller for small sample sets $n<d^{2}/\varepsilon$.
[14] introduced a variant of the projection mechanism, based on the Johnson-Lindestrauss transform ([10]) matching these bounds up to constant factors. For datasets with points of small average Euclidean norm, [5] designed a pure differentially private algorithm achieving a Frobenius norm error of $\tilde{O}(d^{3/4}\sqrt{\operatorname{Tr}\Sigma/n}+\sqrt{d}/n),$ thus going beyond [14] for the special case of matrices satisfying $\operatorname{Tr}\Sigma\ll O(1/\varepsilon)\cdot\min\left\{1/n,\,d^{-1/4}\right\}.$
@@ 1.1 Results
Efficient and streamlined samplers for the nuclear-Laplace law are easy to construct and can be seen as a specific instantiation of the $K$-norm mechanism of [8]. Therefore the algorithm runs in polynomial time.
@@
Theorem 1.1 guarantees bounded error simultaneously for all Schatten norms, including the nuclear norm ($p=1$) and the spectral norm ($p=\infty$), implying spectral discrepancies between the true covariance and the pure differentially private estimate are small under any natural reweighing. For large datasets $n\geqslant d^{2}/\varepsilon$ –typical of real-world applications– the Theorem recovers the guarantees of [14] for the Frobenius norm error. Furthermore, as for $p\geqslant 2$
@@
the error of Theorem 1.1 is optimal for $p\geqslant 2$ in the large sample regime $n\geqslant d^{2}/\varepsilon$. Indeed, the information-theoretic lower bound ([14]) shows that if $n\geqslant d^{2}/\varepsilon$, no pure DP estimator can achieve error $o(d^{3/2}/\varepsilon n)$ in Frobenius norm, and hence no estimator can achieve error $o(d^{1+1/p}/\varepsilon n)$ in Schatten $p$-norm for $p\geqslant 2$. In particular, our spectral norm bound $O(d/\varepsilon n)$ is information-theoretically optimal in the large sample regime $n\geqslant d^{2}/\varepsilon$.
@@
Theorem 1.2 always recovers the result of [14], but improves over it whenever $\operatorname{Tr}\Sigma<d^{2}/n.$ As values of the trace of the covariance matrix can be significantly smaller than $1$, Theorem 1.2 can lead to important accuracy improvements both in the small sample regime $n\leqslant d^{2}/\varepsilon$ where the improvement is of the order $O(1/\operatorname{Tr}\Sigma)$, and in the large sample regime. Theorem 1.2 also strictly improves over the result of [5] (that also captured the dependence on $\operatorname{Tr}(\Sigma)$) in all regimes by a $d^{1/4}$ factor.
@@ 2 Techniques
To illustrate our ideas we start by reviewing the algorithm of [14]. On a high level, this consists of three steps: (1) project the data onto a low dimensional space, (2) add noise via a $K$-norm mechanism, and (3) project the result back onto the space of covariance matrices (the intersection of the nuclear unit ball and the positive semidefinite cone). As the cost of traditional perturbation mechanisms tends to be proportional to the ambient dimension, the goal of the first step is to reduce this privacy cost by embedding the dataset into a small subspace, thus paying a cost proportional only to the effective dimension of the data. The result $\tilde{\Sigma}$ of steps (1) and (2) can then be shown to be private and close to (the projection of) $\Sigma$ up to a Frobenius distance of order $O(d^{3/2}/n\varepsilon)$. The third step is an orthogonal projection onto a convex set containing the ground truth, thus it cannot increase the $\ell_{2}$-distance to $\Sigma$, but, remarkably, for small datasets can provably shrink it to $O(\sqrt{d/n\varepsilon}).$
Despite its utility in terms of the Frobenius norm error, this algorithm cannot be expected to be accurate in other Schatten norms, in a strong sense. The dimensionality reduction step rely on a Johnson-Lindestrauss transform, but such mappings are known not to exist for spaces that are far from being Euclidean and, more specifically, for $p$-norm space with $p\in[1,\infty]\setminus\{2\}$ ([2, 12, 9]).
@@
To bypass this inherent limitation, we do not attempt to explicitly embed the dataset in a low-dimensional space and instead carefully craft the additive noise to inject. Since adjacent covariance matrices $\Sigma,\Sigma^{\prime}$ satisfy $\lVert\Sigma-\Sigma^{\prime}\rVert_{*}\leqslant 2/n$, the $K$-norm mechanism ([8]) with density $p(Z)$ proportional to $\exp(-\varepsilon n\lVert Z\rVert_{*}/2)$ is $\varepsilon$-DP. Note that it is possible to sample from such distribution efficiently, leveraging any efficient weak separation oracle for the nuclear norm [^3] See Lemma A.2. in [8] for more details.. Hence it only remains to bound Schatten norms of $Z$ sampled from this distribution, and in fact, it is enough to bound $\lVert Z\rVert$ by $O(\frac{d}{\varepsilon n})$ since $\lVert Z\rVert_{\textnormal{S}_{p}}\leqslant\lVert Z\rVert\cdot d^{1/p}$.
@@
as [^4] This factorization is standard, and can be found, in particular, in Lemma 1.5.3 from [3].
@@
This implies that to show the desired concentration of the spectral norm, it is enough to study the product of two independent random variables: $R$ and $\max_{i\in[d]}w_{i}$. It is well-known that the distribution of $R$ is [^5] See, for example, Remark 4.2 in [8]. $\mathrm{Gamma}(d^{2},\rho)$, and it is well-concentrated around $\rho d^{2}$ (recall that $\rho=2/(\varepsilon n)$). The result would then follow if we could show that with high probability $\max_{i\in[d]}w_{i}\leqslant 1/d+\tilde{O}(1/d^{3/2})$ as by union bound $\left\lVert Z\right\rVert\leqslant O(d^{2}/\varepsilon\cdot n)\cdot O(1/d)\leqslant O(\tfrac{d}{\varepsilon n}).$
@@
Furthermore, a standard strong convexity argument (e.g. the one used in Lemma 5.2 in [4]) implies that if $r\geqslant\operatorname{Tr}\Sigma$, then the orthogonal projection $\Pi(\Sigma+Z)$ satisfies
@@
where in the before-last step we used the fact that the spectral norm of $Z$ concentrates around $\Theta(d/n\varepsilon)$ with high probability, and in the last step the radius of the nuclear norm ball. The naive choice $r=1$ would recover the result of [14]. However, this choice can be wasteful. We instead use an $\varepsilon$-DP Laplace estimator of $\operatorname{Tr}\Sigma$. Note that with probability at least $1-\beta$, $r\leqslant O\left(\operatorname{Tr}(\Sigma)+\log(1/\beta)/(n\varepsilon)\right)$, and we get the bound as in Theorem 1.2.
@@ 3 Future work
It is natural to ask whether a generalization of both Theorem 1.1 and Theorem 1.2 is possible. Concretely, it would be interesting to determine if there is an estimator satisfying
@@
In addition, it would be interesting to see complementary lower bounds. In particular, currently even in the large sample regime $n\geqslant d^{2}/\varepsilon$ it is only known that our result is optimal for $p\geqslant 2$, but it is not known for $p<2$. And in the small-sample regime, the picture is even less clear.
@@ 4 Analysis of the additive mechanism
(Remark 4.2 in [8]) Let $Z$ have density $p_{\rho}(Z)\propto\exp\left(-\lVert Z\rVert_{*}/\rho\right)$. Then
@@
for a constant $C_{d}>0$, where $d\sigma$ is Lebesgue measure on $\mathbb{R}^{d}$ restricted to $\sigma_{i}\geqslant 0$ and $d\mu_{\mathrm{Haar}}$ denotes Haar probability measure on $\mathrm{O}(d)$ (see Lemma 1.5.3 from [3]).
@@
Since $w$ is strongly log-concave, for each LL–Lipschitz $F$ and each $t>0$,